# Spin Response of a Magnetic Monopole and Quantum Hall Response in Topological Lattice Models through Local Invariants and Light

### v2

@@ 3 Correspondences on a 2D Topological Lattice Model
The symbol $\zeta$ will be fixed such that $\zeta=+1$ at the $K$ Dirac point and $\zeta=-1$ at the $K^{\prime}$ Dirac point. The parameter $M$ is also called a Semenoff mass term related to the Dirac equation [29].
Figure 4: (Top) Honeycomb lattice and Brillouin zone. The rhomboid Brillouin zone is equivalent to the hexagonal one. Within the correspondence for $t_{2}=0.1$ in units of $t$, we introduce the unit vector ${\bf n}_{d}=\frac{{\bf d}}{|{\bf d}|}$ in Eq. (3.1). When the $\bf d$} vector covers the entire sphere then this corresponds to a topological number equal to one e.g. for $M=0$. In this case, Dirac points are located at different poles on the sphere. Within the trivial phase $M=2M_{c}$, favoring the occupancy on one sublattice, both Dirac points then map onto the same pole. In that case, the magnetic field (Berry curvature) does not wrap around the origin completely. This leads to a trivial winding number. (Bottom) Path with fixed azimuthal angle $\varphi$ with the parameter $M=0$ and $t_{2}=0.1$.
@@
For completeness, we illustrate the map from the Brillouin zone, equivalently drawn as a rhomboid, onto the sphere in Fig. 4 that also clarifies the quantum phase transition from a simple geometrical view, i.e. as an equivalent geometry on the sphere encircling or not the origin of the Brillouin zone. Within the topological phase the $K$ Dirac point in red is at north pole at $\theta=0$ and the $K^{\prime}$ Dirac point is at south pole at $\theta=\pi$. Close to the Dirac points, for the analytical evaluations it is useful to introduce [4]
@@
with $v_{F}=\frac{3}{2\hbar}ta$ the Fermi velocity of graphene [43]. We also have the correspondence between the dressed polar angle $\tilde{\theta}$ on the sphere and the deviation from each Dirac point through the wave-vector $|{\bf p}|$ :
@@
where $\zeta$ is $\pm$ at the $K$ and $K^{\prime}$ Dirac points respectively. The formula (35) is very efficient numerically and reproduces well the quantum phase transition at $M=M_{c}=3\sqrt{3}t_{2}$ through the jump of the topological invariant locally resolved at the Dirac points. As illustrated in Fig. 5, this formula works well even when the Berry curvatures vanish at $K$ and $K^{\prime}$.
Figure 5: Berry curvatures for different values of $M$ and $t_{2}$. (Top Left) This corresponds to the physical situation of graphene with a charge density wave substrate leading to Eq. (38) [30]. (Top Right) We are within the topological phase with $C=1$ and the numbers agree with Eq. (38) and also with numerical results in the literature [44]. (Bottom Left) Results within the non-topological phase with $C=0$. (Bottom Right) When increasing $t_{2}$ within the topological phase, the Berry curvature may vanish at $K$ and $K^{\prime}$. Interestingly the formula $C=A_{\varphi}({\bf K}^{\prime})-A_{\varphi}({\bf K})=+1$ from the Dirac points yet works well.
@@ 3.2 Populations and Pseudo-Spin Response as a Local Marker in Momentum Space
It is also possible to re-interpret the sgnsgn function at each Dirac point through a mass sign [47]. If we relate the population in momentum space with the photoluminescence intensity in Ref. [31], then our local analysis may be resolved through the Stokes parameters. In that article, they also introduce the valley Chern number that may be related with our local definition of the topological invariant.
@@ 3.3 Real-Space Analysis and Comparison with the Effective Magnetic Moment of the Monopole
In Fig. 7, we analyse the properties of this local marker in real space comparatively to $\kappa$ in Eq. (20). When varying the parameter $M$, corresponding e.g. to describe the effects of a charge density wave substrate [30], we do reveal a linear response behavior for the pseudo-spin locally on the lattice until the phase transition similar to the analysis of the response of the monopole. The first derivative in yellow of the pseudo-spin is also fixed until the topological phase transition. We report a similar behavior as the susceptibility $\chi=B\frac{\partial\kappa}{\partial M}$ in Eq. (24) when mapping the first Brillouin zone onto the sphere. The first derivative from the real space analysis on the lattice in Fig. 7 is around $\sim 0.33$..., which is approximately half of $\chi$. The susceptibility in yellow in Fig. 7 clearly distinguishes the position of the transition. For completeness, in Fig. 7, we also show in red the second derivative of $\langle\sigma_{z}({\bf R}_{i})\rangle$ as a function of $\frac{M}{M_{c}}$.
@@
The comparison between the local marker in real space $\langle\sigma_{z}({\bf R}_{i})\rangle$ and the theory of the monopole’s spin response is interesting since in general for any observable a careful correspondence between quantities evaluated on the lattice and on the sphere is required. For the quantum Hall response it is possible to re-write the global topological invariant in the plane as an evaluation of the Berry curvature on the sphere through $F_{\theta\varphi}$ with a measure of the area integration as $d\tilde{\theta}d\varphi$ [4]. There is also a simple understanding of this mapping from the Kubo formula for the quantum Hall response [4]. For the pseudo-spin response, this is evaluated on the spherical surface and we have verified that making a flat-space simplification does not reproduce the same results in particular at the transition. The Jacobian of a transformation necessitates some thoughts in general. As long as Eqs. (3.1) are satisfied i.e. with an energy spectrum of the form $\pm\sqrt{(\hbar v_{F}|{\bf p}|)^{2}+(\zeta 3\sqrt{3}t_{2}+M)^{2}}$ around each Dirac point we can justify the results as follows. On the lattice, we evaluate
$\frac{1}{(2\pi)^{2}}\iint dk_{x}dk_{y}\langle\sigma_{z}({\bf k})\rangle.$ (46)
[…]
$\frac{1}{(2\pi)^{2}}\iint dk_{x}dk_{y}\langle\sigma_{z}({\bf k})\rangle\sim\frac{M_{c}^{2}}{\pi}\kappa\sim\frac{1}{\pi}\frac{2}{3}\frac{M}{M_{c}}\sim 0.21\frac{M}{M_{c}}.$ (47)
The precise boundary conditions on the rhomboid Brillouin zone are expected to modify slightly the result. The simple point we are making here is that the Dirac approximation for graphene included through the $d_{x}$ and $d_{y}$ components reveals the linear response behavior of the pseudo-spin in real space. On the lattice, when $t_{2}\sim 0.2t$ the bands start to flatten around the Dirac points [48] such that this approximation becomes more subtle. For larger values of $t_{2}$, numerically we observe some curvature effects on the plateau. Yet, the topological marker introduced on the sphere in Fig. 1 remains applicable for all values of $B$ or $t_{2}$ on the lattice.
@@ 3.4 Relation to Local Responses from Circularly Polarized Light
For $M=0$, one of us have shown that this function at the two Dirac points is precisely related to the square of the topological invariant [6, 4]. Close to $\tilde{\theta}=0$ we can equivalently write
@@
As long as we are within the topological phase, the points $\tilde{\theta}=0$ and $\tilde{\theta}=\pi$ correspond respectively to the points $\theta=0$ and $\theta=\pi$. The functions $\sin^{4}\frac{\tilde{\theta}}{2}$ and $\cos^{4}\frac{\tilde{\theta}}{2}$ can also be interpreted in terms of the topological responses $(A_{\varphi}^{\prime}(\theta_{c}^{-}))^{2}$ and $(A_{\varphi}^{\prime}(\theta_{c}^{+}))^{2}$ ; see Appendix A. The transition probabilities at $\theta=0$ from the right-handed light source ($\Gamma_{+}$) and at $\theta=\pi$ from the left-handed light source ($\Gamma_{-}$), reveal quantized responses $\Gamma_{+}(0)=\Gamma_{-}(\pi)$. The quantized light information at resonance then reveals the definition of the topological invariant resolved at the Dirac points. Fixing $\tilde{\theta}$, when varying $\omega$ we can then resolve the form factor associated to each light response. When crossing the phase transition, the two poles become equivalent from the eigenstates structure. In that case, $\Gamma_{+}(0)=\Gamma_{+}(\pi)$ from the energy conservation. The light responses also reveal the pseudo-spin state at the two Dirac points.
@@
This implies that at the $K^{\prime}$ Dirac point, there is a mirror definition of light polarizations compared to the sphere definitions $\Gamma_{+}(\pi)\rightarrow\Gamma_{LP}({\bf K}^{\prime})$ and $\Gamma_{-}(\pi)\rightarrow\Gamma_{RP}({\bf K}^{\prime})$ where $LP$ and $RP$ designate circular light and right polarizations in the plane. Then, we verify the topological structure ${\Gamma}_{RP}({\bf K})={\Gamma}_{RP}({\bf K}^{\prime})$ within the topological phase and ${\Gamma}_{RP}({\bf K})={\Gamma}_{LP}({\bf K}^{\prime})$ within the trivial phase. For these responses, we have $A_{0}=\frac{\mathcal{E}}{\Delta({\bf k})}$ where $\Delta({\bf k})$ corresponds to the resonance energy gap for a wave-vector $\bf k$} and $\mathcal{E}$ is the electric field. At the transition, the gap is closing at the $K^{\prime}$ point which means effectively that ${\Gamma}_{RP}({\bf K}^{\prime})={\Gamma}_{LP}({\bf K}^{\prime})=0$ if we assume a light source with $\omega\neq 0$.
@@
In this formula the lowest and upper energy eigenstates are introduced as $|l_{\bf k}\rangle$ and $|u_{\bf k}\rangle$, $\lambda$ refers to a right or left circular polarization corresponding to $\lambda=\pm$ respectively. In Sec. 3.1, $p_{x}$ and $p_{y}$ are introduced as wave-vector components measured from each Dirac point. These responses are in accordance with the transition probabilities introduced on the sphere and the selection rules mentioned above. It is then useful e.g. to integrate the responses in frequency. In experiments, for $t_{2}\sim 0.1t$, this is equivalent to fix $\hbar\omega=\Delta({\bf K})$ and integrate on momenta (see Fig. 8) and similarly for the response at the other Dirac point obtained when fixing the resonance $\hbar\omega=\Delta({\bf K}^{\prime})$. We can then verify the precise result
@@
We summarize the results in Fig. 8. Within the topological phase, the circular drive induces two peaks at $K$ and $K^{\prime}$ of height $18\pi\approx 56,55$ if we set $t_{2}=0.1$, $t=1=\mathcal{E}$ and similarly for the lattice spacing. We emphasize on the fact that the heights of these peaks remain identical within the topological phase. When crossing the phase transition, we see clearly that the $K^{\prime}$ Dirac point now interacts with the left-handed circular drive and the peak height remains the same. From the peak heights resolved at the two Dirac points it is now possible to read if the phase is topological or not. It also encodes quantitative information on local Berry curvature and pseudo-spin response.
@@ 4.1 Alternating $\mathbb{Z}$ and $\mathbb{Z}_{2}$ Topological States in Coupled Planes Models
The symbol $\mathcal{R}$} refers to the Ramanujan way of performing the infinite series. We are then wondering what is the physical meaning of this $\frac{1}{2}$ related to the material i.e. how should we think about the thermodynamical limit then ? From the physical point of view, one may argue that an interface between a QAH state and the vacuum or a non-trivial material allows for an interface with a jump of the topological invariant corresponding within our formulation to a situation with $M=M_{c}$ implying a $\frac{1}{2}$ topological number of one massive Dirac point characterized through a $\pi$ Berry phase [24, 33]. In addition, surface states of three-dimensional topological insulators admit a physical situation in a cube with one Dirac point on the upper or bottom surface, that is also described through a similar QAH state [18, 51, 52].
We propose then to unify these two ways of thinking through the infinite alternating series related to observables such as the quantum Hall response and the response to circularly polarized light.
@@
Suppose we apply an electric field $\mathcal{E}$} in each plane where a plane also corresponds to a sphere. In this way we measure the quantum Hall conductivity. Then, the factor $(1-\epsilon j)$ is equivalent to say that the electric field in each plane is effectively ${\mathcal{E}}(1-\epsilon j)$. When $j\rightarrow+\infty$ there exists a limit for which the effective electric field is zero or equivalently the pumped transverse charge (Hall current) is zero. Equivalently, the charge remains at the north pole on each sphere associated to each plane on the right of the interface. The charge cannot be negative because in the view of Newton in Appendix A when reaching the north pole a particle stays there. Therefore, when we meet the plane such that the quantum Hall current is zero then above this limit this is similar as if we insert an infinity of layers which are topologically trivial (e.g. an ordinary insulator of variable width) corresponding then to add many pairs of (11)(1-1) in the Ramanujan series. This implies the presence of an interface that we can model through a topological phase transition in real space through the parameter $M$ : for planes below the interface each plane may be described with a Semenoff mass $M=0$ (or any $M$ smaller than $M_{c}$), then at the interface $M=M_{c}$ and above the interface the material is trivial i.e. $M>M_{c}$. In this sense, we re-interpret the Ramanujan series which is equivalent to
${\mathcal{S}}=1-{\mathcal{S}},$ (70)
@@
${\mathcal{S}}_{1}=\left(1-1+1-1+{\color[rgb]{0,0,1}\frac{1}{2}}\right)+(1-1)+(1-1)+(1-1)+.$... (71)
${\mathcal{S}}_{2}=\left(1-1+1-1+1-{\color[rgb]{0,0,1}\frac{1}{2}}\right)+(1-1)+(1-1)+(1-1)+....={\mathcal{S}}_{1}.$ (72)
In the first summation, the gap closes on the layer 5 at the $K^{\prime}$ Dirac point corresponding to a topological invariant $C=+\frac{1}{2}$ within our analysis. This is precisely equivalent to add a Semenoff mass on the fifth layer such that $M=M_{c}$. The plane 5 is similar to the first plane with $j=0$ and has a topological invariant +1+1 with the definition of the series. The trivial region then may be thought of as planes with $M>M_{c}$. In the second summation, the gap closes on the layer 6 at the $K$ Dirac point corresponding then to a topological invariant $C=-\frac{1}{2}$. In this way, when saying ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ the infinity is indeed reached when we meet the interface because the situations with a even and a odd number of planes in the topological region become identical. Multiplying ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ by $\frac{e^{2}}{h}$ then this is also equivalent to measure the quantum Hall response of the material. This situation can also describe the situation where a material would have a different height in different regions of the sample, i.e. on one region the system will be physically described through the series ${\mathcal{S}}_{1}$ and on one region the system will be physically described through the series ${\mathcal{S}}_{2}$.
@@
Below, we show how the resummation of infinite series can also measure the response to circularly polarized light in the material. The planes associated to an index $j$ which is even will interact with the right-handed circularly polarized light and the planes described with $j$ odd will interact with the left-handed circularly polarized light. If we measure $\kappa^{\prime}$ in Eq. (61) for each plane i.e. the topological invariant, then the total response for the situation of the right-handed circularly polarized light will be
@@
The factor $\frac{1}{j^{\frac{s}{2}}}$ refers to the light propagation in each plane as a power-law attenuation factor for the in-coming vector potential in Sec. 3.4. This series can be re-summed as
@@
with $\zeta(s)$ the Riemann-Zeta function. We suppose that each plane is characterized through the same energy gap at the Dirac points such that light can reach a resonance with each plane. The interesting surprise then is that if we extrapolate $s\rightarrow 0$ i.e. light can reach the infinity in the vertical direction then the result of the measure will be

### v3

@@ 3 Correspondences on a 2D Topological Lattice Model
The symbol $\zeta$ will be fixed such that $\zeta=+1$ at the $K$ Dirac point and $\zeta=-1$ at the $K^{\prime}$ Dirac point. The parameter $M$, favoring the occupancy of a particle (electron) on one sublattice, is called a Semenoff mass term related to the Dirac equation [29].
Figure 4: (Top) Honeycomb lattice and Brillouin zone. The rhomboid Brillouin zone is equivalent to the hexagonal one. Within the correspondence for $t_{2}=0.1$ in units of $t$, we introduce the unit vector ${\bf n}_{d}=\frac{{\bf d}}{|{\bf d}|}$ in Eq. (3.1). When the $\bf d$} vector covers the entire sphere then this corresponds to a topological number equal to one e.g. for $M=0$. In this case, Dirac points are located at different poles on the sphere. Within the trivial phase $M=2M_{c}$, both Dirac points then map onto the same pole. In that case, the magnetic field (Berry curvature) does not wrap around the origin completely. This leads to a trivial winding number. (Bottom) Path with fixed azimuthal angle $\varphi$ with the parameter $M=0$ and $t_{2}=0.1$.
@@
For completeness, we illustrate the map from the Brillouin zone, equivalently drawn as a rhomboid, onto the sphere in Fig. 4 that also clarifies the quantum phase transition from a simple geometrical view, i.e. as an equivalent geometry on the sphere encircling or not the origin of the Brillouin zone. Within the topological phase the $K$ Dirac point in red is at north pole at $\theta=0$ and the $K^{\prime}$ Dirac point in red is at south pole at $\theta=\pi$. Close to the Dirac points, for the analytical evaluations it is useful to introduce [4]
@@
with $v_{F}=\frac{3}{2\hbar}ta$ the Fermi velocity of graphene [44]. We also have the correspondence between the dressed polar angle $\tilde{\theta}$ on the sphere and the deviation from each Dirac point through the wavevector $\bf p$} :
@@
where $\zeta$ is $\pm$ at the $K$ and $K^{\prime}$ Dirac points respectively. The formula (54) is very efficient numerically and it reproduces well the quantum phase transition at $M=M_{c}=3\sqrt{3}t_{2}$ through the jump of the topological invariant locally resolved at the Dirac points. As illustrated in Fig. 5, this formula works well even when the Berry curvatures become very small at $K$ and $K^{\prime}$.
At the topological quantum phase transition, the gap is closing at one Dirac point and the half topological invariant can be seen as a $\pi$ Berry phase or $\pi$ winding number related to the massive Dirac point. From the discussion on the sphere in Sec. 2, this formula $C=A_{\varphi}({\bf K}^{\prime})-A_{\varphi}({\bf K})$ is gauge invariant. We emphasize here that the situation of half topological numbers can be then generalized to a region of the parameters space (e.g. a line) associated to the quantum anomalous Hall semimetal [24, 34, 35]. This formalism builds a correspondence with the physics of surface states of three-dimensional insulators [18, 45, 46] which also reveal a massive topological Dirac point with a $\pi$ winding number or half topological invariant (see also Section 4.6 in Ref. [4]). We mention here recent efforts to observe a half-quantized Hall conductance in semimagnetic topological insulator bilayers [47] related to the $C=1/2$ parity anomaly [7]. A recent work also reports a half-quantized chiral edge current [48].
Figure 5: Berry curvatures for different values of $M$ and $t_{2}$. (Top Left) This corresponds to the physical situation of graphene with a charge density wave substrate leading to Eq. (57) [30]. (Top Right) We are within the topological phase with $C=1$ and the numbers agree with Eq. (57) and also with numerical results in the literature [51]. (Bottom Left) Results within the non-topological phase with $C=0$. (Bottom Right) When increasing $t_{2}$ within the topological phase, there is a deviation from the Dirac approximation for the $d_{x}$ and $d_{y}$ components and the local Berry curvatures become (very) small. Interestingly the formula $C=A_{\varphi}({\bf K}^{\prime})-A_{\varphi}({\bf K})=+1$ from the Dirac points which is exact yet works well.
@@ 3.2 Populations and Pseudo-Spin Response as a Local Marker in Momentum Space
It is also possible to re-interpret the sgnsgn function at each Dirac point through a mass sign [52]. If we relate the population in momentum space with the photoluminescence intensity in Ref. [31], then our local analysis may be resolved through the Stokes parameters. In that article, the authors then introduce the valley Chern number [7].
@@ 3.3 Real-Space Analysis and Comparison with the Effective Magnetic Moment of the Monopole
In Fig. 7, we analyse the properties of this local marker in real space comparatively to $\kappa$ in Eq. (21). When varying the parameter $M$, corresponding e.g. to describe the effects of a charge density wave substrate [30], we do reveal a linear response behavior for the pseudo-spin locally on the lattice until the phase transition similar to the analysis of the response of the monopole. The first derivative in yellow of the pseudo-spin is also fixed until the topological phase transition. We report a similar behavior as the susceptibility $\chi=B\frac{\partial\kappa}{\partial M}$ in Eq. (33). The first derivative from the real space analysis on the lattice in Fig. 7 is around $\sim 0.33$..., which is approximately half of $\chi$. The susceptibility in yellow in Fig. 7 clearly distinguishes the position (location) of the transition. In Fig. 7, we show in red the second derivative of $\langle\sigma_{z}({\bf R}_{i})\rangle$ as a function of $\frac{M}{M_{c}}$.
@@
The comparison between the local marker in real space $\langle\sigma_{z}({\bf R}_{i})\rangle$ and the theory of the monopole’s spin response is interesting since in general for any observable a careful correspondence between quantities evaluated on the lattice and on the sphere is required. For the quantum Hall response it is possible to re-write the global topological invariant in the plane as an evaluation of the Berry curvature on the sphere through $F_{\theta\varphi}$ with a measure of the area integration as $d\theta d\varphi$ [4]. There is also a simple understanding of this mapping from the Kubo formula for the quantum Hall response [4]. For the pseudo-spin response, this is evaluated on the spherical surface and we have verified that making a flat-space simplification does not reproduce the same results in particular at the transition. The Jacobian of a transformation necessitates some thoughts in general. As long as Eqs. (3.1) are satisfied i.e. with an energy spectrum of the form $\pm\sqrt{(\hbar v_{F}|{\bf p}|)^{2}+(\zeta 3\sqrt{3}t_{2}+M)^{2}}$ around each Dirac point we can justify the results as follows. On the lattice, we evaluate
$\frac{1}{(2\pi)^{2}}\iint\langle\sigma_{z}({\bf k})\rangle dk_{x}dk_{y}.$ (65)
In the vicinity of the north pole or of the $K$ Dirac point, from Eq. (52), we can develop $dk_{x}dk_{y}\rightarrow(2\pi)|{\bf p}|d|{\bf p}|$ such that $(\hbar v_{F})^{2}d|{\bf p}||{\bf p}|=d\theta\sin\theta M_{c}^{2}$ with $M_{c}=3\sqrt{3}t_{2}$. We develop the area of the rhomboid Brillouin zone from a disk centered around the $K$ Dirac point. In this correspondence the radius of the sphere is $M_{c}$ whereas the radius of the disk around the $K$ Dirac point is $\hbar v_{F}$. To reproduce the lattice result from the sphere which shows a linear relation between $\theta$ and $|{\bf p}|$ around north pole then this requires e.g. to rescale $\hbar v_{F}=1$ and $M_{c}=1$ corresponding to $t_{2}\sim 0.19$ and $t=\frac{2}{3}$ when setting the lattice spacing to unity. In this way,
$\frac{1}{(2\pi)^{2}}\iint\langle\sigma_{z}({\bf k})\rangle dk_{x}dk_{y}\sim\frac{1}{\pi}\kappa\sim 0.21\frac{M}{M_{c}}.$ (66)
This approximation tends to provide a good justification of why $\langle\sigma_{z}({\bf R}_{i})\rangle$ is related to $\kappa$ and therefore also shows a plateau within the topological phase, as observed numerically in Fig. 7. For larger values of $t_{2}$, numerically we observe some curvature effects on the plateau for $\langle\sigma_{z}({\bf R}_{i})\rangle$ which can be associated to deviations from the parabolic band approximation and to the flattening of the bands [43]. The topological phase transition remains visible. Yet, the topological marker introduced on the sphere in Fig. 1 remains applicable for all values of $B$ or $t_{2}$ on the lattice.
@@ 3.4 Relation to Local Responses from Circularly Polarized Light
For $M=0$, one of us has shown that this function at the two Dirac points is precisely related to the square of the topological invariant [6, 4]. Close to $\tilde{\theta}=0$ we can equivalently write
@@
As long as we are within the topological phase, the points $\tilde{\theta}=0$ and $\tilde{\theta}=\pi$ correspond respectively to the points $\theta=0$ and $\theta=\pi$. The functions $\sin^{4}\frac{\tilde{\theta}}{2}$ and $\cos^{4}\frac{\tilde{\theta}}{2}$ can also be interpreted in terms of the topological responses $(A_{\varphi}^{\prime}(\theta_{c}^{-}))^{2}$ and $(A_{\varphi}^{\prime}(\theta_{c}^{+}))^{2}$ ; see Appendix A. The transition probabilities at $\theta=0$ from the right-handed light source ($\Gamma_{+}$) and at $\theta=\pi$ from the left-handed light source ($\Gamma_{-}$) reveal quantized responses $\Gamma_{+}(0)=\Gamma_{-}(\pi)$. The quantized light information at resonance then reveals the definition of the topological invariant resolved at the Dirac points. Fixing $\tilde{\theta}$, when varying $\omega$ we can then resolve the form factor associated to each light response. When crossing the phase transition, the two poles become equivalent from the eigenstates structure. In that case, $\Gamma_{+}(0)=\Gamma_{+}(\pi)$ from the energy conservation. The light responses also reveal the pseudo-spin state at the two Dirac points.
@@
Compared to the definitions on the sphere, this implies $\Gamma_{+}(\pi)\rightarrow\Gamma_{LP}({\bf K}^{\prime})$ and $\Gamma_{-}(\pi)\rightarrow\Gamma_{RP}({\bf K}^{\prime})$ where $LP$ and $RP$ designate circular light and right polarizations in the plane. Then, we verify the topological structure ${\Gamma}_{RP}({\bf K})={\Gamma}_{RP}({\bf K}^{\prime})$ within the topological phase and ${\Gamma}_{RP}({\bf K})={\Gamma}_{LP}({\bf K}^{\prime})$ within the trivial phase. For these responses, we have $A_{0}=\frac{\mathcal{E}}{\Delta({\bf k})}$ where $\Delta({\bf k})$ corresponds to the resonance energy gap for a wavevector $\bf k$} and $\mathcal{E}$ is the electric field. At the transition, the gap is closing at the $K^{\prime}$ point which means effectively that ${\Gamma}_{RP}({\bf K}^{\prime})={\Gamma}_{LP}({\bf K}^{\prime})=0$ if we assume a light source with $\omega\neq 0$.
@@
In this formula the lowest and upper energy eigenstates are introduced as $|l_{\bf k}\rangle$ and $|u_{\bf k}\rangle$, $\lambda$ refers to a right or left circular polarization corresponding to $\lambda=\pm$ respectively. In Sec. 3.1, $p_{x}$ and $p_{y}$ are introduced as wavevector components measured from each Dirac point. These responses are in accordance with the transition probabilities introduced on the sphere and the selection rules mentioned above. It is then useful e.g. to integrate the responses in frequency. In experiments, for $t_{2}\sim 0.1t$, this is equivalent to fix $\hbar\omega=\Delta({\bf K})$ and integrate on momenta (see Fig. 8) and similarly for the response at the other Dirac point obtained when fixing the resonance $\hbar\omega=\Delta({\bf K}^{\prime})$. We can then verify the precise result
@@
We summarize the results in Fig. 8. Within the topological phase, the right-handed circular drive induces two peaks at $K$ and $K^{\prime}$ of height $18\pi\approx 56,55$ if we set $t_{2}=0.1$, $t=1=\mathcal{E}$ and similarly for the lattice spacing. We emphasize on the fact that the heights of these peaks remain identical within the topological phase. When crossing the phase transition, we see clearly that the $K^{\prime}$ Dirac point now interacts with the left-handed circular drive and the peak height remains the same. From the peak heights resolved at the two Dirac points it is now possible to read if the phase is topological or not. It also encodes quantitative information on local Berry curvature and pseudo-spin response.
@@ 4.1 Alternating $\mathbb{Z}$ and $\mathbb{Z}_{2}$ Topological States in Coupled Planes Models
The symbol $\mathcal{R}$} refers to the Ramanujan way of performing the infinite series. We are then wondering what is the physical meaning of this $\frac{1}{2}$ related to the material i.e. how should we think about the thermodynamical limit then ? From the physical point of view, one may argue that an interface between a QAH state and the vacuum or a non-trivial material allows for an interface with a jump of the topological invariant corresponding within our formulation to a situation with $M=M_{c}$ implying a $\frac{1}{2}$ topological number of one massive Dirac point characterized through a $\pi$ Berry phase [24, 33]. In addition, surface states of three-dimensional topological insulators with one Dirac point on the upper or bottom surface are also described through a similar QAH state [18, 45, 46]. Such surfaces then give rise to a massive Dirac point with a half-quantized quantum Hall conductance which is measured in layered systems [47]. We propose then to unify these two ways of thinking through the infinite alternating series related to observables such as the quantum Hall response and the response to circularly polarized light. It is relevant to emphasize here the recent interest in relating Ramanujan infinite series with topological quantum Hall physics [54].
@@
Suppose we apply an electric field $\mathcal{E}$} in each plane where a plane also corresponds to a sphere. In this way we measure the quantum Hall conductivity. Then, the factor $(1-\epsilon j)$ is equivalent to say that the electric field in each plane is effectively ${\mathcal{E}}(1-\epsilon j)$. When $j\rightarrow+\infty$ there exists a limit for which the effective electric field is zero or equivalently the pumped transverse charge (Hall current) is zero. Equivalently, the charge remains at the north pole on each sphere associated to each plane on the right of the interface. The charge cannot be negative because in the view of Newton in Appendix A when reaching the north pole a particle stays there. Therefore, when we meet the plane such that the quantum Hall current is zero then above this limit this is similar as if we insert an infinity of layers which are topologically trivial (e.g. an ordinary insulator of variable width) corresponding then to add an infinite number of pairs of (11)(1-1) in the Ramanujan series. This implies the presence of an interface that we can model through a topological phase transition in real space through the parameter $M$ introduced in preceding Sections : for planes below the interface each plane may be described with a Semenoff mass $M=0$ (or any $M$ smaller than $M_{c}$), then at the interface $M=M_{c}$ and above the interface the material is trivial i.e. $M>M_{c}$. In this sense, we re-interpret the Ramanujan series which is equivalent to
${\mathcal{S}}=1-{\mathcal{S}}=\frac{1}{2}$ (89)
@@
${\mathcal{S}}_{1}=\left(1-1+1-1+{\color[rgb]{0,0,1}\frac{1}{2}}\right)+(1-1)+(1-1)+(1-1)+....=\frac{1}{2}$ (90)
${\mathcal{S}}_{2}=\left(1-1+1-1+1-{\color[rgb]{0,0,1}\frac{1}{2}}\right)+(1-1)+(1-1)+(1-1)+....={\mathcal{S}}_{1}=\frac{1}{2}.$ (91)
In the first summation, the gap closes on the layer 5 at the $K^{\prime}$ Dirac point corresponding to a topological invariant $C=+\frac{1}{2}$ (mentioned in blue in the equation above) within our analysis. This is precisely equivalent to add a Semenoff mass on the fifth layer such that $M=M_{c}$. The plane 5 is similar to the first plane with $j=0$ and has a topological invariant +1+1 with the definition of the series. The trivial region then may be thought of as planes with $M>M_{c}$. In the second summation, the gap closes on the layer 6 at the $K$ Dirac point corresponding then to a topological invariant $C=-\frac{1}{2}$. In this way, when saying ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ the infinity is indeed reached when we meet the interface because the situations with a even and a odd number of planes in the topological region become identical. Multiplying ${\mathcal{S}}_{1}={\mathcal{S}}_{2}$ by $\frac{e^{2}}{h}$ then this is also equivalent to measure the quantum Hall response of the material. This situation can also describe the situation where a material would have a different height in different regions of the sample, i.e. on one region the system will be physically described through the series ${\mathcal{S}}_{1}$ and on one region the system will be physically described through the series ${\mathcal{S}}_{2}$.
@@
Below, we show how the (physical) regularization of infinite series can also measure the response to circularly polarized light in the material. First, we evaluate the response to circularly polarized light for the mathematical situation of an infinite number of planes with alternating topological numbers. We will then show that this response is equivalent to the materials associated to Eqs. (90) and (91). The planes associated to an index $j$ which is even will interact with the right-handed circularly polarized light and the planes described with $j$ odd will interact with the left-handed circularly polarized light. If we measure $\kappa^{\prime}$ in Eq. (80) for each plane i.e. the topological invariant, then the total response for the situation of the right-handed circularly polarized light will be
@@
The factor $\frac{1}{j^{\frac{s}{2}}}$ refers to the light propagation in each plane through a power-law attenuation factor for the in-coming vector potential in Sec. 3.4. This series can be re-summed as
@@
with $\zeta(s)$ the Riemann-Zeta function. We suppose that each plane is characterized through the same energy gap at the Dirac points such that light can reach a resonance with each plane. The interesting surprise then is that if we extrapolate $s\rightarrow 0$ i.e. light can reach the infinity in the vertical direction such that the result of the measure will be